Consider a firm Theta that spends $60 to produce goods in year 1. In year 2 it sells these goods for $100, but its customers pay their bills with a delay, therefore the payment is not received until year 3. What's the amount of cost of goods sold for each year?

Multiple Choice

  • year 1: $0; year 2: $60; year 3: $0

  • year 1: $0; year 2: $0; year 3: $60

  • year 1: $0; year 2: $100; year 3: $0

  • year 1: $60; year 2: $0; year 3: $0
Answer #1

Cost of goods sold = (starting inventory+ Purchase -closing inventory)

Year 1= (0+60-60)= 0

Year 2= (60+0-0)= 60

Year 3= (0+0-0)=0

Correct answer will be Option (A) Year 1:$0, Year 2:$60, Year 3:$0